Circular No. 8380

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SCORPIUS H. Yamaoka, Kyushu University, reports that Akira Takao, Kitakyu-

shu, Japan, has discovered a possible nova (mag 7.4) on unfiltered CCD images taken with a 120-mm telephoto lens on Aug. 3.583 UT. The new object is located at α = 17^{h}29^{m}18^{s}, δ = −31^{o}46′.0 (equinox 2000.0; uncertainty ∼ 5′′). Nothing was visible on his previous images (limiting magnitudes: July 18.637, [11.9; 30.569, [10.8). An ASAS-3 image taken on Aug. 2.071 also shows this new object at V = 9.9, though the previous ASAS-3 images show nothing at this position (most recently on July 26.271). No known x-ray or bright infrared source is catalogued at this position.

COMETS C/1996 L2, M3; C/2004 L3–L11, M1–M7 (SOHO) Continuation to IAUC 8367 (BZ = B. Zhou):

Comet 1996 UT α_{2000} δ_{2000} Inst. F MPEC
C/1996 L2 June 13.432 5^{h}25.^{m}5 +20^{o}06′ C3 XL 2004-P15
C/1996 M3  20.612 5 50.8 +21 11 C3 XL 2004-P15
Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 L3 June 5.076 4^{h}53.^{m}1 +20^{o}49′ C2  RK 2004-O04
C/2004 L4  7.916 5 04.2 +21 04 C2  XZ 2004-O04
C/2004 L5  8.185 5 04.8 +21 03 C2  BZ 2004-O04
C/2004 L6  9.588 5 10.4 +21 23 C2  XZ 2004-O04
C/2004 L7  9.596 5 08.8 +20 19 C3/2 RK 2004-O04
C/2004 L8  10.554 5 13.3 +20 27 C3/2 KC 2004-O05
C/2004 L9  10.976 5 15.9 +21 16 C2  XZ 2004-O05
C/2004 L10  13.671 5 22.1 +23 16 C2  RK 2004-O05
C/2004 L11  15.338 5 30.9 +21 48 C2  TH 2004-P12
C/2004 M1  18.033 5 42.8 +21 32 C3/2 RK 2004-P12
C/2004 M2  24.888 6 09.1 +20 33 C3/2 JS 2004-P12
C/2004 M3 26.392 6 15.9 +21 40 C2 DC 2004-P12
C/2004 M4  26.967 6 17.6 +21 41 C3/2 TH 2004-P13
C/2004 M5  27.558 6 20.6 +21 45 C2  TH 2004-P13
C/2004 M6  27.558 6 20.6 +21 44 C2  XL 2004-P13
C/2004 M7  27.883 6 21.2 +21 54 C2  TH 2004-P13

All of the above comets are Kreutz sungrazers except for C/2004 L10 (which is a Kracht-group member).

2004 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green